OpenStudy (anonymous):

1. If a and b are any two real numbers, explain how you calculate the difference a – b. Be sure to include the following in your explanation: a. A description of the relationship between addition and subtraction b. How you can use a number line to help with the calculation c. How you can determine the sign of the result

OpenStudy (anonymous):

@Sam_Aka_Sara Ill try

OpenStudy (anonymous):

A - B is found by counting out the quantity A, then removing one for every one in the quantity B. Addition means to get larger by combining quantities. Subtraction means to get smaller by finding how many more there is in one group than another group, then not using the common amount. A number line helps here, since to add is to go to the right, and to subtract is to go to the left. The sign of a subtraction is the sign of the larger of the two numbers involved, if they are opposite signs. If both are the same sign, that is the sign of the answer. Note that if B is negative, this becomes A - (-B) = A + B If both are negative, this is then -A - (-B) = - (A + B) If A is negative, but B positive, you have -A + B -- or effectively B - A